OpenStudy (anonymous):

4x+5y=42 5x+2y=27 using the matrix system

OpenStudy (anonymous):

You will need to know how to modify a matrix using combinations of rows to get the augmented matrix into reduced row eschelon form. Do you know how to do this?

OpenStudy (anonymous):

\[ \left[\begin{matrix}4 & 5 & 42 \\ 5 & 2 & 27\end{matrix}\right] \] => \[ \left[\begin{matrix}1 & 0 & ? \\ 0 & 1 & ?\end{matrix}\right] \]